Numerical Analysis I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

1. Section 4.1: Problems 19, 22, 24, 28, 29.

2. Section 4.2: Problems 5, 10, 12.

3. For section 4.2, problem 12, also show K_{1} = 0 alternatively by assuming the expansion

e = (\frac{2 + h}{2 − h})^{\frac{1}{h}} + C_{1}h^{p_{1}} + C_{2}h^{p_{2}} + ⋯

and find p_{1} numerically.